# Learning to Deblur Polarized Images

### v1

@@ 4.1 Unpolarized image estimator
Overall architecture. We design the first stage as an unpolarized image estimator $g_{1}$, aiming to estimate $\mathbf{I}$ from $\mathbf{B}$. Apparently, the problem handled by $g_{1}$ is similar to deblurring a single conventional image, however, since $\mathbf{B}$ is computed from $\mathbf{B}_{\alpha_{1,2,3,4}}$ which are captured by a polarization camera, the low-level features (e.g., structural and boundary information) would suffer from severe degradation, making the problem more challenging. This is because the DoFP technology adopted by modern polarization cameras induces higher level sensor noise (compared to the same sensor without micro-polarizers) (Kumar et al, 2007; Qiu et al, 2019; Abubakar et al, 2018) so that the noise term $\mathbf{N}$ in Equation (5) and Equation (6) becomes non-negligible. Prominently, we notice that there are two Stokes parameters ($\mathbf{S}_{1,2}$) having similar “appearance” to the gradient distribution $\mathbf{G}$ of the corresponding unpolarized image due to their “differential” nature (i.e., both of them describe the difference between two polarized images according to Equation (4)), as shown in Fig. 2 (left); besides, $\mathbf{S}_{1,2}$ are less affected by motion blur since they often have relatively stronger response near the edges, which means that their blurry counterparts $\hat{\mathbf{S}}_{1,2}$ could retain less-contaminated structural and boundary information. To verify it, we design a simulation experiment on our training set (including 8000 images, see Section 5.2 for details) to quantitatively measure the gap between the sharp/blurry variable pairs, by computing the average value of LPIPS (learned perceptual image patch similarity, which is suitable for measuring the similarity distance in the feature space) (Zhang et al, 2018b) score. Results are shown in Fig. 2 (middle), we can see that the degeneration degree of $\mathbf{S}_{1,2}$ is nearly half of that of $\mathbf{G}$. Therefore, we expect that $\hat{\mathbf{S}}_{1,2}$ can compensate the degraded low-level features, and decide to use them as priors. As shown in the red dashed frame of Fig. 1, this stage can be written as the following equation:
$\mathbf{I}_{\text{guide}}=g_{1}(\mathbf{B},\hat{\mathbf{S}}_{1,2}),$ (7)
@@
Layer details. We first adopt two feature extraction heads (consisting of several convolution layers) $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ to extract features from $\hat{\mathbf{S}}_{1}$ and $\hat{\mathbf{S}}_{2}$ respectively, and adopt a feature fusing block (consisting of a convolution layer and a dense block (Huang et al, 2017)) $\mathcal{F}$ to obtain the fused Stokes features $\mathbf{F}$. Then, we adopt another feature extraction head $\mathcal{H}_{b}$ to extract features from $\mathbf{B}$, and feed them along with $\mathbf{F}$ into a backbone network $\mathcal{B}_{1}$ (a modified autoencoder architecture (Hinton and Salakhutdinov, 2006) similar to the one used in (Zhou et al, 2023b)) to learn the residual between $\mathbf{B}$ and $\mathbf{I}$. For fully utilizing the less-contaminated structural and boundary information retained in $\hat{\mathbf{S}}_{1,2}$, we need to allow $\mathbf{F}$ to directly affect $\mathbf{B}$ across the entire feature domain. Therefore, instead of directly feeding it into the encoder of $\mathcal{B}_{1}$, we propose to use Stokes guided skip connections (SGSC) to substitute the skip connections inside $\mathcal{B}_{1}$ for processing it. SGSC accepts two kinds of inputs: the features from the backbone network and the features from the Stokes parameters. It uses residual bottleneck blocks (He et al, 2016) to filter each feature channel, and uses a squeeze-and-excitation block (Hu et al, 2018) to recalibrate the features by learning normalized weights for recovering fine-grained high-order image structure. Specifically, we adopt a feature downsampling block (consisting of several convolution layers and residual blocks (He et al, 2016)) $\mathcal{D}_{1}$ to downsample $\mathbf{F}$ to get multiscale features of the Stokes parameters ($\mathbf{F}_{1}$ and $\mathbf{F}_{2}$), and adopt two SGSCs ($\mathcal{C}_{1}$ and $\mathcal{C}_{2}$) to process them respectively.
@@ 5.1 Loss function
Content loss. The content loss $L_{c}$ aims to bridge the gap between the restored image $\mathbf{I}_{*}$ (which can be any one of $\mathbf{I}_{\text{guide}}$ and $\mathbf{I}_{\alpha_{1,2,3,4}}$) and its corresponding ground truth for sharper image content, which is defined as
$\begin{split}L_{c}(\mathbf{I}_{*})&=\lambda_{c_{1}}L_{1}(\mathbf{I}_{*})+\lambda_{c_{2}}L_{2}(\mathbf{I}_{*})\\ &+\lambda_{c_{3}}L_{\text{perc}}(\mathbf{I}_{*})+\lambda_{c_{4}}L_{\text{grad}}(\mathbf{I}_{*}),\end{split}$ (10)
where $L_{1}$, $L_{2}$, $L_{\text{perc}}$, and $L_{\text{grad}}$ denote the $\ell_{1}$, $\ell_{2}$, perceptual, and gradient loss respectively, and $\lambda_{c_{1,2,3,4}}$ are set to be 10, 100, 0.1, and 10 respectively. The perceptual loss $L_{\text{perc}}$ is defined as the $\ell_{2}$ loss computed using the feature maps extracted by $VGG_{3,3}$ convolution layer of VGG-19 network (Simonyan and Zisserman, 2014) pretrained on ImageNet (Russakovsky et al, 2015), and the gradient loss $L_{\text{grad}}$ is defined as the $\ell_{2}$ loss computed using the gradients extracted by Laplace kernels.
@@
where $\lambda_{s_{1,2,3,4}}$ are set to be 20, 500, 500, and 500 respectively. Note that the fourth term of $L_{s}$ is adopted to optimize the ratio between $\mathbf{S}_{2}$ and $\mathbf{S}_{1}$ since the calculation of the AoP depends on it (see Equation (3)).
@@ 5.2 Dataset preparation
    Collect polarized images from two public polarized low-light image enhancement datasets (LLCP (Xu et al, 2022) and PLIE (Zhou et al, 2023b)) to serve as the ground truth sharp images.
@@
    Move the corresponding ground truth sharp images along the pixel-wise trajectories to get multiple sharp latent frames.
@@
    Add noise to the blurry images to better simulate the real situation.
@@ 6.1 Evaluation on synthetic data
Since there is no existing approach specially designed for deblurring polarized images, we choose to compare our method to several state-of-the-art methods that solve the closest problems:
  • $\bullet$
    MIMO (Cho et al, 2021) and XYDeblur (Ji et al, 2022) (designed for deblurring a single unpolarized image), which can process the polarized images in a frame-by-frame manner;
  • $\bullet$
    STDAN (Zhang et al, 2022a) (designed for deblurring continuous unpolarized video frames), which can process the polarized images simultaneously in the intensity domain;
  • $\bullet$
    PLIE (Zhou et al, 2023b) (designed for enhancing the polarized low-light images captured by a polarization camera), which can process the polarized images simultaneously in the Stokes domain.
Note that since STDAN (Zhang et al, 2022a) reconstructs the middle frame by referencing 5 neighboring frames, when compared to it we set its input as a video whose frames are stacked by the blurry polarized images in the order of “$[\mathbf{B}_{\alpha_{3}},\mathbf{B}_{\alpha_{4}},\mathbf{B}_{\alpha_{1}},\mathbf{B}_{\alpha_{2}},\mathbf{B}_{\alpha_{3}},\mathbf{B}_{\alpha_{4}},\mathbf{B}_{\alpha_{1}},\mathbf{B}_{\alpha_{2}}]$” to ensure that the output frames are in the order of “$[\mathbf{B}_{\alpha_{1}},\mathbf{B}_{\alpha_{2}},\mathbf{B}_{\alpha_{3}},\mathbf{B}_{\alpha_{4}}]$”.
All of the compared methods are finetuned on our dataset under the supervision of the polarized images $\mathbf{I}_{\alpha_{1,2,3,4}}$. The reason why we should not choose the unpolarized image $\mathbf{I}$ for supervision is because the domain gap between the polarized images and the corresponding unpolarized image is very large, otherwise, the results will be severely corrupted since the unique properties of the polarized images cannot be learned. Besides, for a fair comparison, we choose not to apply additional polarization-related loss functions (e.g., the Stokes loss in Equation (11)) to the compared methods designed for deblurring (MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), and STDAN (Zhang et al, 2022a)) during the training process as other works that aim to enhance polarized images (Hu et al, 2020; Xu et al, 2022; Zhou et al, 2023b) do. This is because we cannot obtain multiple polarized images used for computing the polarization-related parameters in a single forward pass. And if we forcibly change the logic of forward/backward propagation of those methods the results would degenerated.
Visual quality comparisons are shown in Fig. $3$Additional results can be found in the supplementary material.. Note that we not only evaluate the accuracy of the restored DoP $\mathbf{p}$ and AoP $\bm{\theta}$, but also evaluate the quality of the restored unpolarized image $\mathbf{I}$. As for $\mathbf{p}$ and $\bm{\theta}$, our method achieves better performance than other methods, because we explicitly consider to maintain the polarization constraints. Taking the right box of $\mathbf{p}$ in the first group of images as an example, MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), and STDAN (Zhang et al, 2022a) suffer severely from artifacts, and PLIE (Zhou et al, 2023b) tend to generate over-smooth edges. This is because MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), and STDAN (Zhang et al, 2022a) are designed for deblurring unpolarized images so that they can only process the polarized images in a polarization-unaware manner; despite PLIE (Zhou et al, 2023b) is specially designed for processing the polarized images captured by a polarization camera, it aims to handle the low-light enhancement problem, which explores different aspects of polarization properties from the deblurring problem. As for $\mathbf{I}$, our results resemble the ground truth more closely with less distortion. This is because our method can properly deal with the degraded low-level features and less distinctive high-level features.
Figure 3: Qualitative comparisons on synthetic data among our method, MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and PLIE (Zhou et al, 2023b). Quantitative results evaluated using PSNR (P) and SSIM (S) are displayed below each image. The DoP $\mathbf{p}$ and AoP $\bm{\theta}$ are visualized using color maps after normalizing and averaging the RGB channels (as done in (Hu et al, 2020; Zhou et al, 2023b)). Please zoom-in for better details.
@@
Figure 4: Qualitative comparisons on real data among our method, MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and PLIE (Zhou et al, 2023b). The DoP $\mathbf{p}$ and AoP $\bm{\theta}$ are visualized using color maps after normalizing and averaging the RGB channels (as done in (Hu et al, 2020; Zhou et al, 2023b)). Please zoom-in for better details.
Table 1: Quantitative comparisons on synthetic data among our method, MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and PLIE (Zhou et al, 2023b). Bold font indicates the best performance.
Ours 29.12 0.774 17.77 0.330 29.67 0.868
MIMO (Cho et al, 2021) 26.61 0.730 16.51 0.280 25.92 0.830
XYDeblur (Ji et al, 2022) 22.74 0.698 15.83 0.272 26.68 0.850
STDAN (Zhang et al, 2022a) 28.28 0.756 17.06 0.309 27.49 0.862
PLIE (Zhou et al, 2023b) 26.32 0.744 17.24 0.269 28.40 0.853
@@
W/o $\hat{\mathbf{S}}_{1,2}$ 28.41 0.772 17.67 0.329 29.16 0.863
$\mathbf{G}$ instead of $\hat{\mathbf{S}}_{1,2}$ 28.43 0.768 17.70 0.329 29.19 0.864
@@
To evaluate the results quantitatively, we choose to adopt two frequently-used metrics including PSNR and SSIM. Results are shown in Table 1 (also below corresponding examples in Fig. 3). Our model consistently outperforms the compared methods on all metrics.

### v2

@@ 4.1 Unpolarized image estimator
Overall architecture. We design the first stage as an unpolarized image estimator $g_{1}$, aiming to estimate $\mathbf{I}$ from $\mathbf{B}$. Apparently, the problem handled by $g_{1}$ is similar to deblurring a single conventional image, however, since $\mathbf{B}$ is computed from $\mathbf{B}_{\alpha_{1,2,3,4}}$ which are captured by a polarization camera, the low-level features (e.g., structural and boundary information) would suffer from severe degradation, making the problem more challenging. This is because the DoFP technology adopted by modern polarization cameras induces higher level sensor noise (compared to the same sensor without micro-polarizers) (Kumar et al, 2007; Qiu et al, 2019; Abubakar et al, 2018) so that the noise term $\mathbf{N}$ in Equation (5) and Equation (6) becomes non-negligible. Prominently, we notice that there are two Stokes parameters ($\mathbf{S}_{1,2}$) having similar “appearance” to the gradient distribution $\mathbf{G}$ of the corresponding unpolarized image due to their “differential” nature (i.e., both of them describe the difference between two polarized images according to Equation (4)), as shown in Fig. 2 (left); besides, $\mathbf{S}_{1,2}$ are less affected by motion blur since they often have relatively stronger response near the edges, which means that their blurry counterparts $\mathbf{S}^{*}_{1,2}$ could retain less-contaminated structural and boundary information. To verify it, we design a simulation experiment on our training set (including 8000 images, see Section 5.2 for details) to quantitatively measure the gap between the sharp/blurry variable pairs, by computing the average value of LPIPS score (learned perceptual image patch similarity, which is suitable for measuring the similarity distance in the feature space) (Zhang et al, 2018b). Results are shown in Fig. 2 (middle), we can see that the degradation degree of $\mathbf{S}_{1,2}$ is nearly half of that of $\mathbf{G}$. Therefore, we expect that $\mathbf{S}^{*}_{1,2}$ can compensate the degraded low-level features, and decide to use them as priors. As shown in the red dashed frame of Fig. 1, this stage can be written as the following equation:
$\mathbf{I}_{\text{guide}}=g_{1}(\mathbf{B},\mathbf{S}^{*}_{1,2}),$ (7)
@@
Figure 2: Left: Two of the Stokes parameters ($\mathbf{S}_{1,2}$) have similar “appearance” to the gradient distribution $\mathbf{G}$ of the corresponding unpolarized image (please zoom-in for better details). Middle: LPIPS (Zhang et al, 2018b) scores between the blurry/sharp variable pairs, including $\mathbf{S}_{1,2}$ and $\mathbf{G}$. Right: Semantic distance between $\mathbf{I}_{\alpha_{1,2,3,4}}$ and $\mathbf{I}$, $\mathbf{I}_{\alpha_{1,2,3,4}}$ and $\mathbf{B}_{\alpha_{1,2,3,4}}$.
Layer details. We first adopt two feature extraction heads (consisting of several convolution layers) $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ to extract features from $\mathbf{S}^{*}_{1}$ and $\mathbf{S}^{*}_{2}$ respectively, and adopt a feature fusing block (consisting of a convolution layer and a dense block (Huang et al, 2017)) $\mathcal{F}$ to obtain the fused Stokes features $\mathbf{F}$. Then, we adopt another feature extraction head $\mathcal{H}_{b}$ to extract features from $\mathbf{B}$, and feed them along with $\mathbf{F}$ into a backbone network $\mathcal{B}_{1}$ (a modified autoencoder architecture (Hinton and Salakhutdinov, 2006) similar to the one used in (Zhou et al, 2023b)) to learn the residual between $\mathbf{B}$ and $\mathbf{I}$. For fully utilizing the less-contaminated structural and boundary information retained in $\mathbf{S}^{*}_{1,2}$, we need to allow $\mathbf{F}$ to directly affect $\mathbf{B}$ across the entire feature domain. Therefore, instead of directly feeding it into the encoder of $\mathcal{B}_{1}$, we propose to use Stokes guided skip connections (SGSC) to substitute the skip connections inside $\mathcal{B}_{1}$ for processing it. SGSC accepts two kinds of inputs: the features from the backbone network and the features from the Stokes parameters. It uses residual bottleneck blocks (He et al, 2016) to filter each feature channel, and uses a squeeze-and-excitation block (Hu et al, 2018) to recalibrate the features by learning normalized weights for recovering fine-grained high-order image structure. Specifically, we adopt a feature downsampling block (consisting of several convolution layers and residual blocks (He et al, 2016)) $\mathcal{D}_{1}$ to downsample $\mathbf{F}$ to get multiscale features of the Stokes parameters ($\mathbf{F}_{1}$ and $\mathbf{F}_{2}$), and adopt two SGSCs ($\mathcal{C}_{1}$ and $\mathcal{C}_{2}$) to process them respectively.
@@ 5.1 Loss function
Content loss. The content loss $L_{c}$ aims to bridge the gap between the restored image $\mathbf{I}_{\forall}$ (which can be any one of $\mathbf{I}_{\text{guide}}$ and $\mathbf{I}_{\alpha_{1,2,3,4}}$) and its corresponding ground truth for sharper image content, which is defined as
$\begin{split}L_{c}(\mathbf{I}_{\forall})&=\lambda_{c_{1}}L_{1}(\mathbf{I}_{\forall})+\lambda_{c_{2}}L_{2}(\mathbf{I}_{\forall})\\ &+\lambda_{c_{3}}L_{\text{perc}}(\mathbf{I}_{\forall})+\lambda_{c_{4}}L_{\text{edge}}(\mathbf{I}_{\forall}),\end{split}$ (10)
where $L_{1}$, $L_{2}$, $L_{\text{perc}}$, and $L_{\text{edge}}$ denote the $\ell_{1}$, $\ell_{2}$, perceptual, and edge loss respectively, and $\lambda_{c_{1,2,3,4}}$ are set to be 10, 100, 0.1, and 10 respectively. The perceptual loss $L_{\text{perc}}$ is defined as the $\ell_{2}$ loss computed using the feature maps extracted by $VGG_{3,3}$ convolution layer of VGG-19 network (Simonyan and Zisserman, 2014) pretrained on ImageNet (Russakovsky et al, 2015), and the edge loss $L_{\text{edge}}$ is defined as the $\ell_{2}$ loss computed using the edges extracted by Laplace kernels.
@@
where $\lambda_{s_{1,2,3,4}}$ are set to be 20, 500, 500, and 500 respectively. Note that the fourth term of $L_{s}$ is adopted to optimize the ratio between $\mathbf{S}_{2}$ and $\mathbf{S}_{1}$ since the calculation of the AoLP depends on it (see Equation (3)).
@@ 5.2 Dataset preparation
    Collect polarized images from two public polarized low-light image enhancement datasets (LLCP (Xu et al, 2022) and PLIE (Zhou et al, 2023b)) to serve as the ground-truth sharp images.
@@
    Move the corresponding ground-truth sharp images along the pixel-wise trajectories to get multiple sharp latent frames.
@@
    Add noise to the blurry images (as done in other works (Hu et al, 2020; Xu et al, 2022; Zhou et al, 2023b)) to better simulate the real situation.
@@
Table 1: Quantitative comparisons on synthetic data among our method, IPLNet (Hu et al, 2020), ColorPolarNet (Xu et al, 2022), PLIE (Zhou et al, 2023b), MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and MISCFilter (Liu et al, 2024). Bold font indicates the best performance.
Ours 29.12 0.774 17.77 0.330 29.67 0.868
IPLNet (Hu et al, 2020) 19.98 0.702 17.01 0.267 26.08 0.822
ColorPolarNet (Xu et al, 2022) 28.41 0.740 16.99 0.279 25.09 0.812
PLIE (Zhou et al, 2023b) 26.32 0.744 17.24 0.269 28.40 0.853
MIMO (Cho et al, 2021) 26.61 0.730 16.51 0.280 25.92 0.830
XYDeblur (Ji et al, 2022) 22.74 0.698 15.83 0.272 26.68 0.850
STDAN (Zhang et al, 2022a) 28.28 0.756 17.06 0.309 27.49 0.862
MISCFilter (Liu et al, 2024) 28.75 0.763 17.16 0.311 28.45 0.865
Figure 3: Qualitative comparisons on synthetic data among our method, IPLNet (Hu et al, 2020), ColorPolarNet (Xu et al, 2022), PLIE (Zhou et al, 2023b), MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and MISCFilter (Liu et al, 2024). The DoLP $\mathbf{p}$ and AoLP $\bm{\theta}$ are visualized using color maps after normalizing and averaging the RGB channels (as done in (Hu et al, 2020; Zhou et al, 2023b)). Please zoom-in for better details.
@@ 5.3 Training strategy
Table 2: Quantitative comparisons on real data among our method, IPLNet (Hu et al, 2020), ColorPolarNet (Xu et al, 2022), PLIE (Zhou et al, 2023b), MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and MISCFilter (Liu et al, 2024). Bold font indicates the best performance.
CLIP-IQA-$\mathbf{p}$ CLIP-IQA-$\bm{\theta}$ CLIP-IQA-$\mathbf{I}$
Ours 0.506 0.677 0.439
IPLNet (Hu et al, 2020) 0.503 0.628 0.428
ColorPolarNet (Xu et al, 2022) 0.479 0.592 0.366
PLIE (Zhou et al, 2023b) 0.484 0.589 0.274
MIMO (Cho et al, 2021) 0.500 0.667 0.343
XYDeblur (Ji et al, 2022) 0.504 0.624 0.359
STDAN (Zhang et al, 2022a) 0.479 0.593 0.366
MISCFilter (Liu et al, 2024) 0.433 0.582 0.361
@@ 6.1 Evaluation on synthetic data
Since there is no existing approach specially designed for deblurring polarized images, we choose to compare our method to two groups of methods that solve the closest problems. The first group is originally designed for handling the polarized image enhancement tasks, including IPLNet (Hu et al, 2020), ColorPolarNet (Xu et al, 2022), and PLIE (Zhou et al, 2023b). The second group is originally designed for handling the unpolarized image/video deblurring tasks, including MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and MISCFilter (Liu et al, 2024). All methods in the first group can process the polarized images simultaneously in a direct manner. In the second group, STDAN (Zhang et al, 2022a) can process the polarized images simultaneously in an indirect manner since it is designed for deblurring continuous video frames, while MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), and MISCFilter (Liu et al, 2024) can only process the polarized images in a frame-by-frame manner since it is designed for deblurring a single image. Note that since STDAN (Zhang et al, 2022a) reconstructs the middle frame by referencing 5 neighboring frames, when compared to it we set its input as a video whose frames are stacked by the blurry polarized images in the order of “$[\mathbf{B}_{\alpha_{3}},\mathbf{B}_{\alpha_{4}},\mathbf{B}_{\alpha_{1}},\mathbf{B}_{\alpha_{2}},\mathbf{B}_{\alpha_{3}},\mathbf{B}_{\alpha_{4}},\mathbf{B}_{\alpha_{1}},\mathbf{B}_{\alpha_{2}}]$” to ensure that the output frames are in the order of “$[\mathbf{B}_{\alpha_{1}},\mathbf{B}_{\alpha_{2}},\mathbf{B}_{\alpha_{3}},\mathbf{B}_{\alpha_{4}}]$”.
All the compared methods are trained on our dataset (either finetuned or retrained for adaptation to our dataset) under the supervision of the polarized images $\mathbf{I}_{\alpha_{1,2,3,4}}$ (the very same outputs of our method). In practice, we choose to finetune (i.e., train from pre-trained weights) the methods designed for deblurring unpolarized images (Cho et al, 2021; Ji et al, 2022; Liu et al, 2024; Zhang et al, 2022a) and retrain (i.e., train from scratch) the methods designed for enhancing polarized images (Hu et al, 2020; Xu et al, 2022; Zhou et al, 2023b). Such a training strategy for the compared methods could be reasonable and often adopted by other methods (Hu et al, 2020; Xu et al, 2022; Zhou et al, 2023b), since finetuning is commonly used when adapting a model to a specific task or domain that is closely related to the original task the model was trained on (e.g., deblurring unpolarized images vs. deblurring polarized images), while retraining is typically used when the task is different from the original tasks that pre-trained models cannot provide a good starting point (e.g., enhancing low-light polarized images vs. deblurring polarized images). Besides, for a fair comparison, we apply our loss function to the methods designed for enhancing polarized images (Hu et al, 2020; Xu et al, 2022; Zhou et al, 2023b), while preserving the original loss functions for the methods designed for deblurring unpolarized images (Cho et al, 2021; Ji et al, 2022; Liu et al, 2024; Zhang et al, 2022a) during training. This is because our loss function includes a Stokes loss term (Equation (11)), which requires multiple polarized images to compute the polarization-related parameters, while multiple polarized images cannot be obtained within a single forward pass for the methods designed for deblurring unpolarized images. Additionally, forcibly modifying the forward/backward propagation logic of these methods to accommodate our loss function may degrade their performance rather than improve it.
@@
Figure 4: Qualitative comparisons on real data among our method, IPLNet (Hu et al, 2020), ColorPolarNet (Xu et al, 2022), PLIE (Zhou et al, 2023b), MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and MISCFilter (Liu et al, 2024). The DoLP $\mathbf{p}$ and AoLP $\bm{\theta}$ are visualized using color maps after normalizing and averaging the RGB channels (as done in (Hu et al, 2020; Zhou et al, 2023b)). Please zoom-in for better details.
To evaluate the results quantitatively, we choose to adopt two frequently-used metrics including PSNR and SSIM. Results are shown in Table 1. Our model consistently outperforms the compared methods on all metrics. Visual quality comparisons are shown in Fig. $3$Additional results can be found in the supplementary material.. Note that we not only evaluate the accuracy of the restored DoLP $\mathbf{p}$ and AoLP $\bm{\theta}$, but also evaluate the quality of the restored unpolarized image $\mathbf{I}$. As for $\mathbf{p}$ and $\bm{\theta}$, our method achieves better performance than other methods, because we explicitly consider preserving the polarization constraints. Taking the second box of $\mathbf{p}$ as an example, the compared methods either suffer from severe artifacts (e.g., IPLNet (Hu et al, 2020), MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), and STDAN (Zhang et al, 2022a)) or tend to generate over-smooth edges (e.g., ColorPolarNet (Xu et al, 2022), PLIE (Zhou et al, 2023b), and MISCFilter (Liu et al, 2024)). This is because the methods designed for deblurring can only process the polarized images in a polarization-unaware manner, while the methods designed for enhancing polarized images do not take the unique properties of motion blur into consideration. As for $\mathbf{I}$, our results resemble the ground truth more closely with less distortion. This is because our method can properly deal with the degraded low-level features and less distinctive high-level features.
@@
W/o $\mathbf{S}^{*}_{1,2}$ 28.41 0.772 17.67 0.329 29.16 0.863
$\mathbf{G}$ instead of $\mathbf{S}^{*}_{1,2}$ 28.43 0.768 17.70 0.329 29.19 0.864
@@
Figure 5: Results of polarization-based vision applications (including (a) image dehazing (Zhou et al, 2021) and (b) reflection removal (Lyu et al, 2019, 2022)) without deblurring and deblurred by our method and the compared methods (IPLNet (Hu et al, 2020), ColorPolarNet (Xu et al, 2022), PLIE (Zhou et al, 2023b), MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and MISCFilter (Liu et al, 2024)). Please zoom-in for better details.